# FaaF: Facts as a Function for the evaluation of RAG systems

### 06 Mar 2024 v1

@@ 4 Assessment of fact verification formulations in the RAG setting
Facts Formulation (ground truth) Answer Ungrounded Answer Poor Answer
N/A ER F1m N/A ER F1m N/A ER F1m
Prompt(T/F) 0/281 2.1 $*{}^{*}$ 4/281 26.7 77.6 3/281 50.8 61.8
FaaF(T/F) 0/281 1.7 $*{}^{*}$ 0/281 24.9 79.6 0/281 9.9 94.2
FaaF(T/F/N) 0/281 1.4 $*{}^{*}$ 0/281 18.1 85.7 0/281 7.4 95.7
FaaF(T/F)+citation 0/281 0.7 $*{}^{*}$ 0/281 22 82.5 0/281 7.8 95.6
FaaF(T/F/N)+citation 0/281 0.3 $*{}^{*}$ 0/281 23.1 81.7 0/281 7.4 95.8
Table 2: Results on Error Rate (ER) along with F1micro estimated by each fact formulation. F1micro measures accuracy and recall in predicting the false facts. $*{}^{*}$1micro is 0 in the Ground Truth Answer cases because there are no False facts in the ground truth evaluation in this category. T/F indicates True/False and T/F/N are True/False/Not clear from the given passage respectively. N/A indicates the Not answered i.e. the number of facts that the LM did not respond on their validity.
@@ 5 Results
Table 2 presents the non-response rate (N/A), error rate and F1m from the fact verification formulations across the answer categories which are examined.
@@
Prompting for verification is not reliable in cases of incorrect or incomplete information. Although in the case of ground truth answers Prompt(T/F) shows 2.1%ER (in line with what is reported in Min et al. (2023) when retrieval is enabled), we see a sharp rise when we try to verify facts in text from the ungrounded answers (26.7%) and a further deterioration with poor answers (50.8%). Such performance indicates that prompting is not suitable for fact verification in text with unknown information quality and completeness.
FaaF outperforms zero shot prompt in all cases. Formulations which leverage function calling demonstrate a notable improvement in ER and F1micro over all cases (answer type, FaaF configuration). The improvement is most pronounced in the poor answer category where ER is reduced from 50.8% (prompting) to less than 10% with FaaF methods denoting a paradigm shift in LMeval’s ability to identify missing information in the input text. The milder improvement seen in the ungrounded answer category suggests that elucidating the fact verification in the presence of incorrect and indirectly relevant information is more challenging for LMeval. Lastly, in the case of prompting, LMeval failed to respond on 4/281 of the examined facts for the ungrounded answer and 3/281 for the poor answer. Conversely, LMeval returned a valid response for all facts, in all FaaF configurations and answer types.
LMs tend to overestimate fact truthfulness overall. The human evaluation of factual accuracy in ground truth answers is 100% i.e. every fact is True (Table 1). This coincides with the lowest ER scores in Table 2, irrespective of the fact verification approach.
@@
Rejecting facts is more difficult in the presence of tangentially relevant information when using FaaF. It is evident from Table 2 that FaaF approaches struggle more with ungrounded answers than with poor answers. It should be noted that this observation is reversed when using prompt(T/F), where the ER is higher in poor answers, due to the significant overestimation of truthfulness this approach shows in the absence of supporting information. The key difference between the two answer types is that ungrounded answers typically contain incorrect or tangentially relevant information to the facts that are examined whilst poor answers usually lack information altogether. This is because ungrounded answers are generated by an LM without any context thus, depending on the popularity and prevalence of the theme of the question, a degree of factual hallucination can be expected.
Consider the following example from WikiEval:
@@
Figure 3: LMeval call count for a full evaluation of WikiEval. FaaF formulations result in more than five times less LM calls.
@@
Table 3: LMeval token count for the tested fact formulations, for the factual recall evaluation of WikiEval dataset.
Providing a “not clear” option helps. We observe a reduction of the error rate and corresponding increase in F1m when we add the option for LMeval to respond with Not clear from the given passage which is mapped to False as a post-generation step in the parsing function. The helpful mechanism in this instance is that we provide a needed third option to LMeval when the token probability distribution between True / False is not clearly indicating one over the other. Rejecting a statement due to conflicting evidence and due to lack of evidence are both as valid reason as they are distinct to each other. Using False to capture both rejection scenarios has proved to lead to more false positives than providing LMeval with the option to distinguish between them.
Citations help only in the presence of correct and clear information. The positive impact of adding citations is most evident in the ground truth answer category where the provided text always contain the required evidence to support the facts. In this instance, asking for text evidence from LMeval results in avoiding some false negative verifications. The beneficial mechanism is associated with inserting a preliminary step to the fact verification process - of the explicit use of evidence from the input text. This aligns with the findings reported in Wei et al. (2023) regarding the chain-of-thought method.
Interestingly, for the other answer categories, the citation benefit becomes less clear and even reversed. ER is relatively stable in poor answers and it is seen to increase in the case of ungrounded answer. In detail, we notice two conflicting trends: firstly, citations can prevent false positives by highlighting the absence of supporting text for a given fact statement. Conversely, in a few instances, the lack of explicit citations leads to false negatives, as LMeval fails to recognize a fact as true, even though a holistic human interpretation of the answer would deem it accurate.
LMeval usage efficiency
Beyond the improvements in ER and F1m, another advantage in encapsulating multiple facts into a function object is the that of efficiency. Figure 3 shows reduction in the number of calls to LMeval required for the factual evaluation of WikiEval by more than five times. The number of calls to LMeval required for FaaF corresponds to the number of texts to be evaluated, whereas, in prompt-based fact verification, it matches the number of fact statements.
Table 3 indicates that FaaF(T/F) uses half the prompt tokens and less than one-fifth of the completion tokens required by prompt-based fact verification for a complete evaluation on WikiEval(all answer types). Particularly in the setting of RAG systems optimisation, the efficiency and scalability of an evaluation framework is of high importance since it is proportionally related to the number of RAG configurations which can be tested.
Including citations and the option Not clear from the given passage as acceptable responses in FaaF affects the token count, significantly increasing it. However, it remains below the token requirements of prompt-based fact verification. In essence, FaaF makes possible fact verification with citations at a lower token cost than prompt-based verification without citations.
@@ 6 Conclusions & future work
We show that prompt-based fact verification is prone to overestimating the truthfulness of fact statements in texts with inaccurate and/or missing information. In particular the error rate of prompting can reach up to 50% when the text under review is significantly lacking in information. In such challenging situations, presenting the facts as a function to the LM significantly enhances both the error rate and the F1m measure of the verification process. Using FaaF, we observe that texts with tangentially relevant and inaccurate information’s are more likely to cause false positives than texts with missing or incomplete information. By testing various configurations with FaaF, we find that by including a “not clear” option to the True/False dichotomy helps overall. The impact of asking for citations before fact verification is sensitive to the quality and coverage of information in the examined text. While beneficial when fact-supportive information is present or clearly absent, this approach can lead to false negatives if the text can support the fact as a whole but lacks direct, extractable citations.
Finally we report significant cost and time efficiency improvements between prompting and function based fact representations. Generally, using FaaF leads to a reduction in both the number of calls to the LM and the number of tokens needed for fact verification by a multiple-factor.
@@
While the advantages of using function calls for fact verification are significant, there is certain variance expected in the results due to the stochastic nature of LM generation. Further extensive testing is necessary to solidify the results presented, especially probing into the influence of citations in the performance of verification using function objects. Although, the WikiEval dataset has been instrumental in providing a convenient way to compare fact verification performance across various text qualities, it is relatively small, comprising only 50 question/answer pairs.
Additionally, the results shown with FaaF are sensitive to the instructions passed in the function object’s metadata, which highlights the need for additional research into different FaaF configurations. Other open questions include the maximum number of fact statements and the maximum permissible length for a fact that can be incorporated into a function object and whether token count is the sole limitation or if there are performance implications as well.

### 08 Apr 2024 v2

@@ 4 Assessment of fact verification formulations in the RAG setting
(ground truth) Answer Ungrounded Answer Poor Answer
Facts Formulation N/A ER F1m N/A ER F1m N/A ER F1m
gpt-3.5-turbo Prompt(T/F) 0/281 1.4 $*{}^{*}$ 0/281 27.4 76.5 16/281 55.3 56.9
FaaF(T/F) 0/281 1 $*{}^{*}$ 0/281 23.1 81 0/281 17.4 89.6
FaaF(T/F/N) 0/281 1 $*{}^{*}$ 0/281 25.9 78 0/281 18.5 88.8
FaaF(T/F)+citation 0/281 1 $*{}^{*}$ 26/281 28.2 75.3 71/281 15.7 90
FaaF(T/F/N)+citation 0/281 1.4 $*{}^{*}$ 26/281 31.3 71 31/281 23.2 85.5
claude-3-sonnet Prompt(T/F) 0/281 1 $*{}^{*}$ 0/281 42.7 58 2/281 77.7 26.4
FaaF(T/F) 0/281 1.7 $*{}^{*}$ 0/281 28.8 75.9 0/281 14.9 91.1
FaaF(T/F/N) 0/281 2 $*{}^{*}$ 0/281 27.7 76.9 0/281 14.2 91.6
FaaF(T/F)+citation 0/281 0.7 $*{}^{*}$ 0/281 30.6 73.9 0/281 14.5 91.4
FaaF(T/F/N)+citation 0/281 1 $*{}^{*}$ 0/281 27.4 76.7 0/281 14.9 91.3
gpt-4-turbo Prompt(T/F) 0/281 1.7 $*{}^{*}$ 1/281 24.2 80.1 9/281 44.1 68.4
FaaF(T/F) 0/281 1 $*{}^{*}$ 0/281 22.4 82 0/281 8.5 95.1
FaaF(T/F/N) 0/281 1.4 $*{}^{*}$ 0/281 17.7 86.3 0/281 6.7 96.2
FaaF(T/F)+citation 0/281 0.7 $*{}^{*}$ 6/281 15.6 88.2 15/281 7.5 95.8
FaaF(T/F/N)+citation 0/281 1 $*{}^{*}$ 6/281 16 87.7 0/281 9.2 94.8
claude-3-opus Prompt(T/F) 0/281 1.7 $*{}^{*}$ 0/281 42.3 58.8 2/281 76.3 28.7
FaaF(T/F) 0/281 3.2 $*{}^{*}$ 0/281 15.3 88.8 0/281 6.7 96.2
FaaF(T/F/N) 0/281 3.5 $*{}^{*}$ 0/281 14.5 89.3 0/281 4.9 97.2
FaaF(T/F)+citation 0/281 0.3 $*{}^{*}$ 0/281 24.1 80.8 0/281 5 97
FaaF(T/F/N)+citation 0/281 0.7 $*{}^{*}$ 0/281 20.9 83.6 0/281 7.8 95.6
Table 2: Results on Error Rate (ER) along with F1micro estimated by each fact formulation. F1micro measures accuracy and recall in predicting the false facts. $*{}^{*}$1micro is 0 in the Ground Truth Answer cases because there are no False facts in the ground truth evaluation in this category. T/F indicates True/False and T/F/N are True/False/Not clear from the given passage respectively. N/A indicates the Not answered i.e. the number of facts that the LM did not respond on their validity due to due to errors in its response format.
@@ 5 Results
Table 2 presents the non-answer rate (N/A), Error Rate and F1m from the examined fact verification formulations and LMs, across the answer categories which are examined.
@@
Prompting for fact verification is not reliable in cases of incorrect or incomplete information. Although in the case of ground truth answers Prompt(T/F) ER is in the low percentage points in all LMs—in line with what is reported in Min et al. (2023) when retrieval is enabled—we see a sharp rise when we try to verify facts in text from the ungrounded answers and a further deterioration with poor answers with ER exceeding 50% and 70% with GPT and Claude LMs respectively. Although the failure mechanisms are distinct between Claude and GPT, the performance in both cases indicates that prompting coupled with word-matching is not a suitable approach for fact verification in text with unknown information quality and completeness.
In the case of gpt-4-turbo and gpt-3.5-turbo, the high error rates are attributed to an overall overestimation of the factual truthfulness of a fact statement given a reference text which does not support it. Regarding claude-3-opus and claude-3-sonnet, the exceedingly high ER is primarily due to the erroneous parsing of the LM’s response. In detail, on many occasions the verification response from Claude LMs contains both True and False words, which exposes the fragility of word-matching as a means of parsing the LM’s response. Although qualitative evaluation found Claude models to be less prone to overestimation than GPT LMs, the increased verbosity of the responses some times includes phrases like ”To determine if the claim is true or false based on the given passage…”, ”…cannot be determined as true or false.” or ”…given the lack of information in the passage, we cannot determine whether the claim is True or False.” which mislead the interpretation of the response when using word matching.
FaaF outperforms prompting in fact verification accuracy. Formulations which leverage function calling demonstrate a notable improvement in ER and F1micro particularly in the cases of ungrounded and poor answer with all LMs. The improvement is most pronounced in the poor answer category where all models show lower ER scores from 30 to 70 percentage points compared to prompting, denoting a paradigm shift in the model’s ability to identify unsupported facts in the examined text. In addition, the larger LMs (claude-3-opus and gpt-4-turbo) preform notably better in the poor answer category than their respective inferior versions (claude-3-sonnet and gpt-3.5-turbo).
@@
Rejecting facts is more difficult in the presence of tangentially relevant information when using FaaF. It is evident from Table 2 that FaaF approaches struggle more with ungrounded answers than with poor answers. The key difference between the two answer types is that ungrounded answers typically contain incorrect or tangentially relevant information to the facts that are examined whilst poor answers usually lack information altogether. This is because ungrounded answers are generated by an LM without any context thus, depending on the popularity and prevalence of the theme of the question, a degree of factual hallucination can be expected. It is apparent that, from the perspective of fact verification, presence of indirectly relevant (but not useful) information in the generated text is more dangerous than lack of information.
Consider the following example from WikiEvalFacts:
@@
LMs tend to overestimate fact truthfulness overall. The human evaluation of factual accuracy in ground truth answers is 100% i.e. every fact is True (Table 1). This coincides with the lowest ER scores in Table 2, irrespective of the fact verification approach. False positive verifications are responsible almost exclusively for the observed error rates with all language models demonstrating excellent verification performance when the facts can be directly supported from the given text.
Providing a “not clear” option helps the larger LMs. We observe a reduction of the error rate and corresponding increase in F1m in claude-3-opus and gpt-4-turbo, when we include the option for LMeval to respond with Not clear from the given passage which is mapped to False as a post-generation step in the invoked function. The helpful mechanism in this instance appears to be that we provide a needed third option to LMeval when the token probability distribution between True / False is not clearly indicating one over the other. Rejecting a statement due to conflicting evidence and due to lack of evidence are both as valid rejection reasons as they are distinct to each other. Using False to capture both rejection scenarios has proved to lead to more false positives than providing LMeval with the option to distinguish between them. The fact that the improvement is only seen in the more capable LMs supports this view since they are more capable for complex tasks and language comprehension.
Asking for citations helps in the presence of correct and clear information but can also lead to false positives otherwise. The positive impact of adding citations is most evident in the ground truth answer category where the provided text always contain the required evidence to support the facts. In this instance, asking for text evidence from LMeval results in avoiding some false negative verifications. The beneficial mechanism is associated with inserting a preliminary step to the fact verification process—of the explicit use of evidence from the input text. This aligns with the findings reported in Wei et al. (2023) regarding the chain-of-thought method.
Interestingly, for the other answer categories, the citation benefit becomes less clear and even reversed. ER is relatively stable in poor answers but it is seen to increase in the case of ungrounded answer when citation arguments are included in FaaF. In detail, we notice the following conflicting effects: firstly, citations can prevent false positives by highlighting the absence of supporting text for a given fact statement when they are left empty by the LM, which is beneficial. Secondly, in other cases they can cause false positives when they contain an indirectly relevant excerpt or an excerpt which only supports partially the fact in question. Consider the following example:
Ungrounded answer:
The human climate niche refers to the range of climatic conditions in which humans can thrive and maintain a sustainable population. It encompasses various factors such as temperature, …
Fact to verify:
The human climate niche refers to the range of climate conditions that have supported human life and activities over the past thousand years
FaaF(T/F)+citation – claude-3-opus:
LM citation: ”The human climate niche refers to the range of climatic conditions in which humans can thrive and maintain a sustainable population.”
LM response: True
FaaF(T/F) – claude-3-opus:
LM response: False
Human:
Manual annotation: False
In the example above, only part of the fact statement can be supported from the ungrounded answer (i.e. there is no evidence that the human climate niche refers to the past thousand years). By asking for the citation, the LM captures the partially supporting excerpt and concludes that the fact is True (which is a false positive) but when the same LM verifies the fact without citation, it correctly rejects it.
It is the net effect of the above competing behaviours which determines the impact of adding citations to the overall evaluation performance. Further, results show that Claude LMs are more sensitive to the adverse effects of citations compared to the GPT family, as seen in ungrounded answers (Table 2).
Figure 3: LMeval call count for a full evaluation of WikiEvalFacts. FaaF formulations result in more than five times less LM calls considering an average of 5.6 fact statements per QA pair.
@@
claude-3-opus Prompt(T/F) 163,513 102,305 265,818
FaaF(T/F) 113,092 18,495 131,587
FaaF(T/F/N) 124,051 20,242 144,293
FaaF(T/F)+citation 181,993 44,146 226,139
FaaF(T/F/N)+citation 192,952 45,323 238,275
@@
Table 3: LMeval token count for the tested fact formulations, for the factual recall evaluation of WikiEvalFacts dataset.
Claude LMs are more reliable than GPT in correctly formatting the response for function calling. The capacity of the LM to return a correctly formatted response for function calling is distinct to their ability for accurate fact verification. In this work we allowed only one chance to respond correctly (regarding format) for all experiments and we report the cases of an failed responses for all LMs in the N/A columns in Table 2. The ER and F1micro metrics are calculated considering only the correctly formatted LM responses.
As evident in Table 2, the XML format used in Claude LMs results to more reliable response formatting compared to JSON format used by GPT models. Claude LMs returned a correctly formatted response 100% of the time when using FaaF whereas GPT LMs produced some failed attempts. The failure mechanism in GPT models appears to be directly related to the citations mode—all formatting failures are seen in FaaF+citation formulations (Table 2). Looking more closely, we find that when the citation of a fact is null, there is a risk that the LM will return null in the fact verification argument as well (but only True or False are accepted according to the type annotations of the function object definition), which results to failed invocation of the FaaF function object. In the case of gpt-4-turbo, the above failure mechanism is much less pronounced compared to gpt-3.5-turbo, which shows up to 25% failure rate (71/281) in returning a correctly formatted response in the poor answers, where the citations are expected to be null most of the time.
We attempted to include mistral-large LM in our study but the failure rates associated the response formatting was prohibitive (80% in some cases). It is worth noting that the failure mechanisms of mistral-large were different to what we observed in GPT LMs. In detail, on many occasions mistral-large failed to return all the expected arguments in its response and in other cases, in stead of returning the expected JSON with the function arguments, it responded with a long-form text, which included the function arguments as part of the natural language response (which led to failure when parsing).
FaaF requires less than one-fifth of the calls to the LM to perform fact verification compared to prompting. This corresponds to the average number of facts which are examined for each text (answer type) in WikiEvalFacts. Thus, the degree of efficiency improvement using FaaF is proportional to the number of facts we can encapsulate in the function object, there by avoiding their individual verification.
Table 3 presents the token counts for the tested verification formulations and model families. The token counts include tokens used for all the answer types in WikiEvalFacts dataset in each scenario. Only one model from the GPT and Claude families is included since the token count differences between models in the same family are not significant. Lastly, it should be noted that although we can constrain the completion tokens from the ML parameters we chose not to do so, to examine the token usage in the unconstrained scenario.
GPT LMs using JSON format are significantly more efficient than Claude with XML in token usage. Considering the differences between gpt-4-turbo and claude-3-opus, the observed increase in prompt and completion tokens in the FaaF approaches is associated to the tags used in the XML format which is expected from claude function calling (versus the more succinct JSON format that is expected by gpt). In the case of prompt tokens, the one-shot system prompt (which also contains XML tags) used in claude-3-opus further contributes to the token count. A respective improvement in speed is also noted, with GPT LMs being faster than Claude.
Focusing at the completion tokens, the sharp increase (more than 4X) between gpt-4-turbo and claude-3-opus seen in the case of prompt-based verification is attributed to the extra verbosity by claude-3-opus. When prompted to verify a single fact, gpt-4-turbo most times returns small phrase containing True/False. Conversely, claude-3-opus always returns a coherent explanation to justify the verification. It should be noted that the same prompt was used in both models.
FaaF significantly reduces token usage, requiring fewer than one-fifth the completion tokens needed for prompt-based verification. Following the trend seen in the required LM calls, the most notable reduction in token count is achieved when replacing prompt-based verification with FaaF(T/F). In this scenario, with gpt-4-turbo half of the prompt tokens and less than one-fifth of the completion tokens are required and with claude-3-opus, two thirds of the prompt tokens and less than one-sixth of of the completion tokens are required for fact verification for a complete evaluation on WikiEval(all answer types). Including citations and the response option Not clear from the given passage progressively increases the token count due to the additional information we include in the LM verification. However, in most cases, it still remains below the token requirements of prompt-based fact verification.
@@ 6 Conclusions & future work
We show that prompt-based fact verification is prone to overestimating the truthfulness of fact statements in texts with inaccurate and/or missing information. In particular the error rate of prompting can exceed 70% when the text under review is significantly lacking in information. In such challenging situations, presenting the facts as a function (FaaF) to the language model significantly enhances the its ability to verify facts accurately. The improvement comes from leveraging a more structured generation mode of the language model. This is achieved by generating type-annotated function call arguments instead of natural language text. Additionally, it avoids the unreliable method of exact word matching to parse verification responses, a common issue in prompt-based verification methods.
Using FaaF, we observe that texts with tangentially relevant and inaccurate information are more likely to cause false positives than texts with missing or incomplete information. By testing various configurations with FaaF, we find that by including a “not clear” option to the True/False dichotomy helps the larger LMs. The impact of asking for citations before fact verification is sensitive to the quality and coverage of information in the examined text and not beneficial in many cases.
Additionally, we report significant cost and time efficiency improvements between prompting and FaaF fact representations. Generally, using FaaF leads to a reduction in both the number of calls to the LM and the number of tokens needed for fact verification by a multiple-factor.
We examine the latest commercial language models and find significant improvement in fact verification accuracy between the larger LMs (gpt-4-turbo and claude-3-opus) and their smaller and faster counterparts (gpt-3.5-turbo and claude-3-sonnet).
GPT models using JSON function representation show a sensitivity in returning a correctly formatted response for function calling when asking from the LM to return citations to support fact verification. In contrast, Claude models using XML function representation are more reliable in this regard, consistently delivering correctly formatted responses.
Language model claude-3-opus using XML slightly outperformed gpt-4-turbo using JSON in fact verification accuracy using FaaF but required significantly more tokens to achieve that due to the more verbose format of XML tags compared to JSON.
@@
While the advantages of using function calls for fact verification are significant, there is certain variance expected in the results due to the stochastic nature of LM generation. Further extensive testing is necessary to solidify the results presented, especially probing into the influence of using XML or JSON as the means of passing the function to the LM in the performance of verification using function objects. Although, the WikiEval dataset has highlighted the importance of testing fact verification in challenging conditions and provided a convenient way to compare fact verification performance across various text qualities, it is relatively small, comprising only 50 question/answer pairs.
Additionally, the results shown with FaaF are sensitive to the instructions passed in the function object’s metadata, which highlights the need for additional research and optimisation of FaaF configurations and the interplay of the function argument’s metadata. One example of such interplay is the observation that in GPT models using JSON, including arguments for citation in the FaaF object, caused a rise in the type validation errors due to the LM confusing the different type annotations of different arguments. Other open questions include the maximum number of fact statements and the maximum permissible length for a fact that can be incorporated into a function object and whether token count is the sole limitation or if there are performance implications as well.